I am to use interval notation for the solution set:
Please help
3x+9=3(x-3)
Here is another one:
3x<9 and 2x+1>3
Thanks, lackinmath

Found 2 solutions by funmath, stanbon:
Answer by funmath(2933)

Hi lackinmath,
Still haven't changed your nickname, huh?
I am to use interval notation for the solution set:
3x+9=3(x-3)
3x+9=3x-9
3x-3x+9=3x-3x-9
0+9=0-9
9=-9 There is no solution. They usually put this / over top of 0 to indicate that in interval notation.
:
Here is another one:
3x<9 and 2x+1>3
3x/3<9/3 and 2x+1-1>3-1
x<3 and 2x>2
x<3 and 2x/2>2/2
x<3 and x>1
"and" means that x is both less than 3 and greater than 1. X is between 3 and 1.
In interval notation: (1,3)
Happy Calculating!!!!
funmath

Answer by stanbon(75887)

I am to use interval notation for the solution set:
Please help
3x+9=3(x-3)
3x+9=3x-9
9=-9
Answer: the null set
--------------------
Here is another one:
3x<9 and 2x+1>3
x<3 and 2x>2
x<3 and x>2
(-inf,3)and(2,inf)
-----------
Cheers,
Stan H.